A Bibliography of Publications of Bernhard Nebel

Bernhard Nebel University of Ulm

Department of Artificial Intelligence

Oberer Eselsberg

D-89069 Ulm Germany

Tel: +49/731/502-4122 FAX: +49/731/502-4119

E-mail: nebel@ki.informatik.uni-ulm.de (Internet)

30 November 2023 Version 1.11

Abstract

This bibliography records publications of Bernhard Nebel.

Title word cross-reference

+ [BN93a, BN93b, BN95a, BN95b]. PPP [AGH^{+}93].

10 [BMNuHHN79, NP76]. 18th [NDF94].

21st [BHN97].

3rd [NSR92].

94 [NDF94]. 97 [BHN97].

Abstract [AFG^{+}95]. Access [HCM^{+}83a, HCM^{+}83b, HBC^{+}84c]. Action [HKNP91, FKKB^{+}05]. Actions [HKNP92c].

ADL [KNHD97]. advances [BHN97, NDF94]. Algebra [NB93, NB94c, NB95]. Allen [NB93, NB94c, NB95]. am [Neb80]. Analyse [MNPS76]. Analysis [BHN^{+}92, BHN^{+}93, BHN^{+}94, HKNP92b, HKNP92a, HKNP94, Neb89a, Neb89b, NK93b, NK95, PHKN91]. Anatomy [vLNPS87]. Annual [BHN97, NDF94]. ANS [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c, MN83a, MN83b, MMN84, NM82a, NM82b, CvHH^{+}83, Neb82]. Approach [LN07]. Approaches [Neb91c, Neb91b, Neb91a, Neb92b]. Arbeiten [Neb80]. argumentation [DNT02]. Artificial [Neb95a, Neb96, BHN97, NDF94]. Aspects [vLNS88, KN89]. assumption [DNT02]. assumption-based [DNT02]. Attributive [NS89, NS90, NS91a, NS91b]. Automatic [BN94]. axioms [THN05].

BACK [vLNPS85b, vLNPS85a, vLNPS87, Neb87a, Neb88, NvL88, PvLNS87]. Background [HCM^{+}83a, HCM^{+}83b]. Balancing [NvL87a, NvL87b]. Base [HBC^{+}84c, vLNS88, Neb94, Neb95b, SN86a, SN86b, Neb98]. Based [AGH^{+}93, NS85, NS86, Neb91c, Neb91b, DNT02, Neb91a, Neb92b]. Basic [GHHN99, GHH^{+}99]. Belief [Neb89a, Neb89b, Neb91a, Neb91c, Neb91b, Neb92b, Neb98]. Bericht [KN89, MMN84]. Beyond [HCM^{+}83a, HCM^{+}83b]. Bitter [SvLNP86]. Building [BN94].

calculi [RN07]. Calculus [SN01, RN99]. Case [SvLNP86]. class [Neb97]. Classification [PN93]. cognition [FKKB^{+}05]. Combining [PN93]. Commented [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c]. Comparative [PHKN91]. compilability [Neb00]. Compiler [Neb80]. Compilers [NP76, BMNuHHN79]. Complex [BN90, BN94, Neb91d]. Complexity [BN93a, BN93b, BN95a, BN92, BN95b, Neb87a, Neb88, NB91, Neb91d, Neb92a, NB92, NK92, NK93a, Neb94, NB94a, NB94b, Neb95b, SN01, DNT02, RN99]. Complexity-Theoretic [NK92, NK93a]. Computational [BN92, Neb87a, Neb88, NB91, Neb91e, Neb92a, NB92, NB94a, Neb95a, Neb95c, SN01, DNT02, Neb96]. Conference [BHN97, NSR92, NDF94, FKKB^{+}05]. connection [RN99]. Consistency [vLNPS85b]. constraint [RN07]. Cooperating [Neb01]. Cooperative [DGN02]. COPANET [Neb80, Neb80]. COPANET-Compiler [Neb80]. Crossing [SN01]. CS [DGN02, GHHN99, GHH^{+}99, INW02, WAD^{+}01, WKD^{+}02]. Cycles [Neb87b, Neb91e].

Data [BN90, HBC^{+}84c, Neb91d]. Database [BN94]. Datenbankzugang [MN83a, MN83b]. Datenbasen [NM82a, NM82b]. Decidability [SN01]. DECSystem [BMNuHHN79, NP76]. DECSystem-10 [BMNuHHN79, NP76]. Default [Neb91c, Neb91b, DNT02, Neb91a]. defense [THN05]. Description [BLNN95, NS91a, NS91b]. Descriptions [NS89, NS90]. Design [SN86a, SN86b]. development [HCM^{+}83a, HCM^{+}83b]. Dialoging [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c]. Diskursbereichen [CvHH^{+}83]. Diskursbereiches [MN83a, MN83b]. Diverse [HCM^{+}83a, HCM^{+}83b]. does [Neb85a, Neb85b]. Doing [WAD^{+}01]. Domain [HCM^{+}83a, HCM^{+}83b]. Domain-Independence [HCM^{+}83a, HCM^{+}83b]. Double [SN01]. Double-Crossing [SN01]. Dump [NP76]. dynamische [DN83].

efficiency [Neb97]. Efficient [RN01]. Einführung [OKvLN93]. Empirical [BHN^{+}92, BHN^{+}93, BHN^{+}94, HKNP92b, HKNP92a, HKNP94, NK93b, NK95]. Encoding [DNK97]. Equations [BBN^{+}91, BBN^{+}93]. Erweiterung [BMNuHHN79, NP76]. Erzeugung [NP76]. evaluating [Neb97]. Evaluation [INW02]. Experience [HCM^{+}83a, HCM^{+}83b]. Experiment [NS85, NS86]. expressive [Neb00]. Expressiveness [NB94b]. Expressivity [BBN^{+}91, BBN^{+}93]. Extending [KNHD97].

facts [NDK97]. Feature

[BBN^{+}91, BBN^{+}93]. First [PN93, Neb95c]. fit [Neb85a, Neb85b]. Form [SN86a, SN86b]. Formal [KN89]. Formalisms [NS91a, NS91b, Neb00]. Foundations [BN90, LN94b, LN94a]. fragment [RN99]. frame [Neb85a, Neb85b]. Frauenchiemsee [FKKB^{+}05]. Freiburg [BHN97, DGN02, GHH^{+}99, INW02, WAD^{+}01, WKD^{+}02, GHHN99]. Functional [BBN^{+}91, BBN^{+}93].

Generation [NS85, NS86, NK92, NK93a, NK93b, NK95, SN86a, SN86b, NDK97]. German [BHN97, NDF94, HCM^{+}83a, HCM^{+}83b]. Germany [BHN97, FKKB^{+}05, NDF94]. get [BHN^{+}92, BHN^{+}93, BHN^{+}94]. Gets [NS85, NS86]. Global [DGN02]. graphs [KNHD97]. Group [WAD^{+}01]. Guide [LN94b, PvLNS87].

HAM [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c, MN83a, MN83b, MMN84, NM82a, NM82b, CvHH^{+}83, Neb82]. HAM-ANS [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c, MN83a, MN83b, MMN84, NM82a, NM82b, CvHH^{+}83, Neb82]. hard [Neb97, Neb98]. HASY [MNPS76]. heterogene [NM82a, NM82b]. Hierarchical [LN07]. Highly [HCM^{+}83a, HCM^{+}83b]. Horn [Neb97]. Hotel [HBC^{+}84a]. Hybrid [NvL87a, NvL87b, NvL88, Neb89d, Neb90a].

Ignoring [NDK97]. II [HBC^{+}84b]. III [HBC^{+}84c]. im [Neb80]. Incompleteness [vLNPS85b]. Independence [HCM^{+}83a, HCM^{+}83b]. Information [NP91c, NP91b, NB94b, RN98, NP91a]. Inherently [Neb89c, Neb90b]. Inheritance [BN94, Neb91d, PN93]. Integration [GHHN99, GHH^{+}99, HKNP91, NvL87a, NvL87b]. Intelligence [Neb95a, BHN97, NDF94, Neb96]. Intelligenter [NM82a, NM82b]. Intelligenz [Neb95d]. Intellimedia [AFG^{+}95]. interaction [FKKB^{+}05]. International [BLNN95, NPvL91a, NPvL91b, NSR92, FKKB^{+}05]. Interrechner [Neb80]. Interrechner-Kommunikation [Neb80]. Interval [NB93, NB94c, NB95]. Intractable [Neb89c, Neb90b]. irrelevant [NDK97].

Issues [NvL87a, NvL87b]. IV [FKKB^{+}05].

KI [BHN97, CvHH^{+}83, NDF94]. KI-94 [NDF94]. KI-97 [BHN97]. KI-System [CvHH^{+}83]. Know [NS85, NS86]. Knowledge [BBH^{+}91, LN94b, LN94a, vLNS88, NS85, NS86, NvL87a, NvL87b, Neb89a, Neb89b, NSR92, PSOKK^{+}90, SvLNP86, SN86a, SN86b, BBH^{+}90]. Knowledge-Base [vLNS88, SN86a, SN86b]. Knowledge-Based [NS85, NS86]. Kommunikation [Neb80]. Komplexitätsanalysen [Neb95d]. Korrespondenz [MN83a, MN83b]. KR&R [Neb92a]. KRIS [BHN^{+}92, BHN^{+}93, BHN^{+}94]. Künstlichen [Neb95d].

langsame [Neb83a, Neb83b]. Language [HCM^{+}83a, HCM^{+}83b, HBC^{+}84a, HBC^{+}84b, HBC^{+}84c, NS85, NS86, SN86a, SN86b]. Languages [PSOKK^{+}90]. Lesson [Neb01]. Level [Neb89a, Neb89b]. LISP [Neb84, Neb83a, Neb83b, Neb85c, DN83]. Lispmaschinen [Neb84]. Lispmaschinen-LISP [Neb84]. Localization [GHHN99, GHH^{+}99]. Logic [BBH^{+}90, BBH^{+}91, NS85, NS86, DNK97]. Logical [NS85, NS86, SN86a, SN86b]. Logical-Form [SN86a, SN86b]. Logics [BBN^{+}91, BBN^{+}93, BLNN95, HKNP91, HKNP92c, NPvL91a, NPvL91b, Neb92c]. Logik [OKvLN93]. Logisches [Neb85c]. loop [Neb85a, Neb85b].

Making [BHN^{+}92, BHN^{+}93, BHN^{+}94].

Management

[vLNS88, NP91c, NP91b, NP91a].

Managing [NB94b]. Matters [Neb92a]. Maximal [NB93, NB94c, NB95, RN99]. methods [RN01]. Modell [MN83a, MN83b]. Models [BN90, Neb91d]. Modification [NK92, NK93a]. Möglichkeit [NP76]. monotonic [DNK97]. Mortem [NP76]. move [BHN^{+}92, BHN^{+}93, BHN^{+}94]. Multimedia [AFG^{+}95]. Multiple [BN94, Neb91d]. Multirobot [GHHN99, GHH^{+}99].

Natural [HCM^{+}83a, HCM^{+}83b, HBC^{+}84a, HBC^{+}84b, HBC^{+}84c, NS85, NS86, SN86a, SN86b]. natürlicher [MNPS76]. natürlichsprachliche [MN83a, MN83b, NM82a, NM82b]. natürlichsprachlicher [CvHH^{+}83, MN83a, MN83b]. Navigation [SN01]. Negation [BBN^{+}91, BBN^{+}93]. Networks [KN89]. NIGEL [NS85, NS86]. non [DNK97]. non-monotonic [DNK97]. Nonmonotonic [PN93].

Object [BN90, BN94, Neb91d]. October [FKKB^{+}05]. Operations [Neb94, Neb95b]. operators [NDK97]. Optimierung [BMNuHHN79]. Optimization [BHN^{+}92, BHN^{+}93, BHN^{+}94]. ORD [Neb97].

papers [FKKB^{+}05]. Papiere [Neb85c]. Part [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c]. PASCAL [BMNuHHN79, NP76]. PASCAL-Compilers [NP76]. PDDL [THN05]. Performance [INW02]. Personalized [AGH^{+}93]. Perspective [NK92, NK93a, Neb95a, Neb96]. Physical [Neb01]. Pills [SvLNP86]. Plan [AGH^{+}93, NB92, NK92, NK93a, NK93b, NB94a, NK95, NDK97]. Plan-Based [AGH^{+}93]. Planning [BN93a, BN93b, BN92, BN95b, NB94a, BN95a, DNK97, KNHD97, Neb00]. Playing [Neb01]. Polynomial [Neb92a]. Post [NP76]. Post-Mortem-Dump [NP76]. power [Neb00]. Presenter [AGH^{+}93]. Principles [NSR92]. Problems [NB91, DNK97, Neb97]. Proceedings [BLNN95, NSR92, BHN97, NDF94].

Programm [MNPS76]. Programmieren [Neb85c]. programs [DNK97]. Project [AGH^{+}93]. Projection [NB91, NB92, NB94a]. Properties [Neb91e, Neb95c]. Proposal [BBH^{+}90, BBH^{+}91, AGH^{+}93]. propositional [Neb00].

Qualitative [LN07, NB94b, Neb95c, RN07, SN01, Neb97, RN99, RN01]. Quantifizierung [MN83a, MN83b].

RAT [HKNP92c]. Realisierung [Neb80]. Reasoning [LN07, Neb87a, Neb88, NvL88, NS89, Neb89c, NS90, Neb90a, Neb90b, Neb91c, Neb91b, NP91c, NP91b, NSR92, NB93, NB94c, Neb95c, NB95, PN93, RN98, DNT02, FKKB^{+}05, Neb91a, NP91a, Neb97, RN99, RN01, RN07]. region [RN99]. Reise [MMN84]. Related [NB91]. Relations [NB93, NB94c, NB95]. Reliable [GHHN99, GHH^{+}99]. Report [NPvL91b]. Representation [BBH^{+}91, BHN^{+}92, BHN^{+}93, BHN^{+}94, HKNP91, HKNP92b, HKNP92a, HKNP92c, HKNP94, LN94b, LN94a, LN07, NvL87a, NvL87b, NS89, Neb89d, NS90, Neb90a, NSR92, Neb92c, Neb94, Neb95b, PSOKK^{+}90, PHKN91, SvLNP86, BBH^{+}90]. Reservation [HBC^{+}84a]. Rest [NS91a, NS91b]. Results [BN93a, BN93b, BN95b, Neb95c, BN95a]. Reuse [NK93b, NK95]. revise [Neb98]. Revised [FKKB^{+}05]. Revision [Neb89a, Neb89b, Neb90a, Neb91c, Neb91b, Neb94, Neb95b, Neb91a, Neb92b]. Right [WAD^{+}01]. Robotic [GHHN99, GHH^{+}99, Neb01]. Robots [Neb01].

Saarbrucken [NDF94]. SAS

[BN93a, BN93b, BN95a, BN95b]. Scene [HBC^{+}84b]. Schemata [BN94]. Schemes [Neb94, Neb95b]. selected [FKKB^{+}05]. Self [GHHN99, GHH^{+}99]. Self-Localization [GHHN99, GHH^{+}99]. Semantic [KN89]. Semantics [Neb91e, Neb94, Neb95b]. semantisches [MN83a, MN83b]. Seminar [NPvL91b]. Sensing [DGN02]. Sensor [GHHN99, GHH^{+}99]. September [BHN97, NDF94]. Sessions [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c]. Situation [HBC^{+}84a]. Skills [GHHN99, GHH^{+}99]. Soccer [GHHN99, GHH^{+}99, Neb01]. Solving [Neb97]. Some [vLNS88, NB91]. Sort [BBN^{+}91, BBN^{+}93]. Spatial [LN07, Neb95c, RN98, FKKB^{+}05, RN99, RN01, RN07, FKKB^{+}05]. Speicherallokationsstrategie [DN83]. Sprache [MNPS76, Neb83a, Neb83b]. Step [PN93]. Story [BN92]. Study [SvLNP86]. Subclass [NB93, NB94c, NB95]. subset [KNHD97]. Subsumption [PSOKK^{+}90]. Supporting [BN94]. Syntaktische [MN83a, MN83b]. syntaktischen [MNPS76]. Syntax [Neb91c, Neb91b, Neb92b, Neb91a]. Syntax-Based [Neb91c, Neb91b, Neb92b, Neb91a]. System [CvHH^{+}83, HCM^{+}83a, HCM^{+}83b, HBC^{+}84a, HBC^{+}84b, HBC^{+}84c, vLNPS85a, vLNPS87, PvLNS87, NM82a, NM82b, Neb84]. Systemrahmen [Neb82]. Systems [BHN^{+}92, BHN^{+}93, BHN^{+}94, HKNP92b, HKNP92a, HKNP94, HCM^{+}83a, HCM^{+}83b, vLNS88, NvL87a, NvL87b, Neb89d, Neb90a, Neb92c, PHKN91].

Taking [NS85, NS86]. Team [GHHN99, GHH^{+}99]. Techniques [BHN^{+}92, BHN^{+}93, BHN^{+}94]. Temporal [NB91, NB92, NB93, NB94a, NB94b, NB94c, NB95, Neb97]. Term [PSOKK^{+}90]. Terminal [HBC^{+}84a, HBC^{+}84b, HBC^{+}84c]. Terminological [BBH^{+}90, BBH^{+}91, BHN^{+}92, BHN^{+}93, BHN^{+}94, HKNP91, HKNP92b, HKNP92a, HKNP92c, HKNP94, Neb87a, Neb87b, Neb88, Neb89c, Neb90b, NPvL91a, NPvL91b, Neb91e, NP91a, NP91c, NP91b, Neb92c, PHKN91]. Theoretic [NK92, NK93a]. Theoretical [BN90, NK93b, NK95]. Thing [WAD^{+}01]. Time [Neb92a]. Topological [RN98]. Tour [MMN84]. Tractable [NB93, NB94c, NB95, RN99]. Traffic [HBC^{+}84b].

Übersetzungs [Neb84].

Übersetzungs-System [Neb84]. UCI [Neb84, DN83]. UCI-LISP [Neb84, DN83]. ULM [Neb84]. Uncertainty [BBN^{+}91, BBN^{+}93]. Understanding [BN92]. unterschiedlichen [CvHH^{+}83]. USA [MMN84]. USA-Reise [MMN84]. User [PvLNS87]. using [HKNP92c, Neb97, RN07].

Validation [BN94, NB92, NB94a]. vanilla [Neb85a, Neb85b]. verdeckten [Neb80]. versus [NK92, NK93a, NK93b, NK95]. Video [AFG^{+}95]. View [DGN02, NS85, NS86]. Volume [LN94b]. vs [NB94b].

Weiterführende [Neb80]. well [Neb85a, Neb85b]. WIP [AFG^{+}95]. Wissens [NM82a, NM82b]. Wissens- [NM82a, NM82b]. Wissensbasierter [CvHH^{+}83]. Wissensrepräsentation [OKvLN93]. Workshop [BLNN95, NPvL91a, NPvL91b, KN89, Neb85c]. World [NS91a, NS91b].

Zugang [CvHH^{+}83]. Zugriff [NM82a, NM82b]. zum [Neb85c]. zur [MNPS76, NP76].

References

Andre:1995:WMI

[AFG^{+}95] Elisabeth André, Wolfgang Finkler, Winfried Graf, Karin Harbusch, Anne Heinsohn, Jochen, Kilger, Bernhard Nebel, Profitlich Hans-Jürgen, Thomas Rist, Wolfgang Wahlster, Andreas Butz, and Anthony Jameson. WIP: From multimedia to intellimedia (abstract of video). In Proceedings of the 14th International Joint Conference on Artificial Intelligence, pages 2053–2054. Morgan Kaufmann, San Mateo, August 1995.

Andre:1993:PPB

[AGH^{+}93] Elisabeth André, Winfried Graf, Jochen Heinsohn, Bernhard Nebel, Hans-Jürgen Profitlich, Thomas Rist, and Wolfgang Wahlster. PPP: Personalized plan-based presenter — Project Proposal. DFKI Document D-93-05, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, May 1993.

Baader:1990:TKR

[BBH^{+}90] Franz Baader, Hans-Jürgen Bürckert, Jochen Heinsohn, Bernhard Hollunder, Jürgen Müller, Bernhard Nebel, Werner Nutt, and Hans-Jürgen Profitlich. Terminological Knowledge Representation: A proposal for a terminological logic. DFKI Technical Memo TM-90-04, Saarbrücken, Germany, 1990. A revised version has been published in [?].

Baader:1991:TKR

[BBH^{+}91] Franz Baader, Hans-Jürgen Bürckert, Jochen Heinsohn, Bernhard Hollunder, Jürgen Müller, Bernhard Nebel, Werner Nutt, and Hans-Jürgen Profitlich. Terminological knowledge representation: A proposal for a terminological logic. In B. Nebel, K. von Luck, and C. Peltason, editors, International Workshop on Terminological Logics. DFKI, Dagstuhl, Germany, 1991. This paper is a revised version of [BBH^{+}90]. The workshop proceedings have been published as DFKI Document D-91-03, as KIT Report 89 and as IWBS Report 184.

Baader:1991:EFL

[BBN^{+}91] Franz Baader, Hans-Jürgen Bürckert, Bernhard Nebel, Werner Nutt, and Gert Smolka. On the expressivity of feature logics with negation, functional uncertainty, and sort equations. DFKI Research Report RR-91-01, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, January 1991. Revised version published in Journal of Logic, Language, and Information.

Baader:1993:EFL

[BBN^{+}93] Franz Baader, Hans-Jürgen Bürckert, Bernhard Nebel, Werner Nutt, and Gert Smolka. On the expressivity of feature logics with negation, functional uncertainty, and sort equations. Journal of Logic, Language, and Information, 2(1):1–18, 1993. CODEN JLLIEN. ISSN 0925-8531. URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/baader-
et-al-jlli93.ps.gz.

Baader:1992:EAO

[BHN^{+}92] Franz Baader, Bernhard Hollunder, Bernhard Nebel, Hans-Jürgen Profitlich, and Enrico Franconi. An empirical analysis of optimization techniques for terminological representation systems or “making KRIS get a move on”. In B. Nebel, W. Swartout, and C. Rich, editors, Principles of Knowledge Representation and Reasoning: Proceedings of the 3rd International Conference, pages 270–281. Morgan Kaufmann, San Mateo, October 1992.

Baader:1993:EAO

[BHN^{+}93] Franz Baader, Bernhard Hollunder, Bernhard Nebel, Hans-Jürgen Profitlich, and Enrico Franconi. An empirical analysis of optimization techniques for terminological representation systems or “making KRIS get a move on”. DFKI Research Report RR-93-03, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, February 1993. Revised version published in Applied Intelligence.

Baader:1994:EAO

[BHN^{+}94] Franz Baader, Bernhard Hollunder, Bernhard Nebel, Hans-Jürgen Profitlich, and Enrico Franconi. An empirical analysis of optimization techniques for terminological representation systems or “making KRIS get a move on”. Applied Intelligence, 4(2):109–132, May 1994. CODEN APITE4. ISSN 0924-669X. URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/baader-
etal-applied-intelligence-
94.ps.

Brewka:1997:KAA

[BHN97] Gerhard Brewka, Christopher Habel, and Bernhard Nebel, editors. KI-97, advances in artificial intelligence: 21st Annual German Conference on Artificial Intelligence, Freiburg, Germany, September 9–12, 1997: proceedings, volume 1303 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CO-DEN LNCSD9. ISBN 3-540-63493-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G466 1997. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1303.htm;   http:
//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=1303.

Borgida:1995:PIW

[BLNN95] Alexander Borgida, Maurizio Lenzerini, Daniele Nardi, and Bernhard Nebel. Proceedings of the international workshop on description logics. Rapporto 07.95, Universita degli Studi di Roma “La Sapienza”, Dipartimento di Informatica e Sistemica, Rome, Italy,

1995. URL ftp://disparcs.
dis.uniroma1.it/pub/DL-
95/Proceedings.ps.gz.

Bruegge:1979:OED

[BMNuHHN79] Bernd Brügge, Klaus Mühle, Bernhard Nebel, and Helmut Faasch und Hans-Hellmut Nagel. Optimierung und Erweiterung des DECSystem-10 PASCAL Compilers. Mitteilung IfI-HH-M65/79, Institut für Informatik, Universität Hamburg, Hamburg, Germany, February 1979.

Bergamaschi:1990:TFC

[BN90] Sonia Bergamaschi and Bernhard Nebel. Theoretical foundations of complex object data models. Technical Report 74, CIOC-CNR, Bologna, Italy, December 1990.

Baeckstroem:1992:CCP

[BN92] Christer Bäckström and Bernhard Nebel. On the computational complexity of planning and story understanding. In B. Neumann, editor, Proceedings of the 10th European Conference on Artificial Intelligence, pages 349–353. Wiley, New York, NY, USA, August 1992.

Backstrom:1993:CRSa

[BN93a] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. Research Report LiTH-IDA-R-93-34, Department of Computer and Information Science, Linköping University, Linköping, Sweden, 1993. A revised version of this paper has been published in Computational Intelligence.

Backstrom:1993:CRSb

[BN93b] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. In Proceedings of the 13th International Joint Conference on Artificial Intelligence, pages 1430–1435. Morgan Kaufmann, San Mateo, August 1993. A long version of this paper has been published in Computational Intelligence.

Bergamaschi:1994:ABV

[BN94] Sonia Bergamaschi and Bernhard Nebel. Automatic building and validation of complex object database schemata supporting multiple inheritance. Applied Intelligence, 4(2):185–204, May 1994. CODEN APITE4. ISSN 0924-669X. URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/bergamaschi-
nebel-applied-intelligence-
94.ps.

Backstrom:1995:CRS

[BN95a] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. Comput. Intelligence, 11(4): 625–655, 1995. ISSN 0824-7935. Baeckstroem:1995:CRS

[BN95b] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. Computational Intelligence, 11(4):625–655, 1995. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic). URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/backstrom-
nebel-ci-95.ps.

Christaller:1983:WNZ

[CvHH^{+}83] Thomas Christaller, Walther von Hahn, Wolfgang Hoeppner, Heinz Marburger, Katharina Morik, Bernhard Nebel, and Wolfgang Wahlster. Wissensbasierter natürlichsprachlicher Zugang zu unterschiedlichen Diskursbereichen mit dem KI-System HAM-ANS. Bericht ANS-12, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, 1983.

Dietl:2002:CFG

[DGN02] Markus Dietl, Jens-Steffen Gutmann, and Bernhard Nebel. CS freiburg: Global view by cooperative sensing. Lecture Notes in Computer Science, 2377:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2377/23770133.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2377/23770133.
pdf.

Dannenberg:1983:DSF

[DN83] Rolf Dannenberg and Bernhard Nebel. Eine dynamische Speicherallokationsstrategie für UCI-LISP. Memo ANS-20, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, November 1983.

Dimopoulos:1997:EPP

[DNK97] Yannis Dimopoulos, Bernhard Nebel, and Jana Koehler. Encoding planning problems in non-monotonic logic programs. Lecture Notes in Computer Science, 1348: 169–181, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimopoulos:2002:CCA

[DNT02] Yannis Dimopoulos, Bernhard Nebel, and Francesca Toni. On the computational complexity of assumptionbased argumentation for default reasoning. Artificial Intelligence, 141(1-2):57–78, 2002. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Freksa:2005:SCI

[FKKB^{+}05] Christian Freksa, Markus Knauff, Bernd Krieg-Brückner, Bernhard Nebel, and Thomas Barkowsky, editors. Spatial cognition IV: reasoning, action, interaction: international conference on spatial cognition 2004, Frauenchiemsee, Germany, October 11– 13, 2004. Revised selected papers, volume 3343 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CO-DEN LNCSD9. ISBN 3-540-25048-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387 .S74 2004. URL http:

//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=3343;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b106616.

Gutmann:1999:CFR

[GHH^{+}99] Jens-Steffen Gutmann, Wolfgang Hatzack, Immanuel Herrmann, Bernhard Nebel, Frank Rittinger, Augustinus Topor, Thilo Weigel, and Bruno Welsch. The CS freiburg robotic soccer team: Reliable selflocalization, multirobot sensor integration, and basic soccer skills. Lecture Notes in Computer Science, 1604:93– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1604/16040093.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1604/16040093.
pdf.

Gulmann:1999:CFR

[GHHN99] J.-S. Gulmann, W. Hatzack, I. Herrmann, and B. Nebel. The CS Freiburg robotic soccer team: Reliable selflocalization, multirobot sensor integration, and basic soccer skills. Lecture Notes in Computer Science, 1604:93–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoeppner:1984:DHAa

[HBC^{+}84a] Wolfgang Hoeppner, Stephan Busemann, Thomas Christaller, Heinz Marburger, Katharina Morik, and Bernhard Nebel. Dialoging HAM-ANS: Commented terminal sessions with a natural language system — Part I: Hotel reservation situation. Memo ANS-23, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, 1984.

Hoeppner:1984:DHAb

[HBC^{+}84b] Wolfgang Hoeppner, Stephan Busemann, Thomas Christaller, Heinz Marburger, Katharina Morik, and Bernhard Nebel. Dialoging HAM-ANS: Commented terminal sessions with a natural language system — Part II: Traffic scene. Memo ANS-23, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, 1984.

Hoeppner:1984:DHAc

[HBC^{+}84c] Wolfgang Hoeppner, Stephan Busemann, Thomas Christaller, Heinz Marburger, Katharina Morik, and Bernhard Nebel. Dialoging HAM-ANS: commented terminal sessions with a natural language system — Part III: Data base access. Memo ANS-23, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, 1984.

Hoeppner:1983:BDIa

[HCM^{+}83a] Wolfgang Hoeppner, Thomas Christaller, Heinz Marburger, Katharina Morik, Bernhard Nebel, Michael O’Leary, and Wolfgang Wahlster. Beyond domain-independence: Experience with the development of a German natural language access system to highly diverse background systems. In Proceedings of the 8th International Joint Conference on Artificial Intelligence, pages 115–121. William Kaufmann, Karlsruhe, Germany, August 1983.

Hoeppner:1983:BDIb

[HCM^{+}83b] Wolfgang Hoeppner, Thomas Christaller, Heinz Marburger, Katharina Morik, Bernhard Nebel, Michael O’Leary, and Wolfgang Wahlster. Beyond domain-independence: Experience with the development of a German natural language access system to highly diverse background systems. Bericht ANS-16, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, April 1983.

Heinsohn:1991:IAR

[HKNP91] Jochen Heinsohn, Daniel Kudenko, Bernhard Nebel, and Hans-Jürgen Profitlich. Integration of action representation in terminological logics. In C. Peltason, K. von Luck, and C. Kindermann, editors, Terminological Logic Users Workshop — Proceedings; KIT Report 95, page 117. KIT Group, Department of Computer Science, TU Berlin, Berlin, Germany, December 1991.

Heinsohn:1992:EATb

[HKNP92a] Jochen Heinsohn, Daniel Kudenko, Bernhard Nebel, and Hans-Jürgen Profitlich. An empirical analysis of terminological representation systems. In Proceedings of the 10th National Conference of the American Association for Artificial Intelligence, pages 767–773. MIT Press, Cambridge, MA, USA, July 1992. An extended and revised version of this paper has been published in Artificial Intelligence.

Heinsohn:1992:EATa

[HKNP92b] Jochen Heinsohn, Daniel Kudenko, Bernhard Nebel, and Hans-Jürgen Profitlich. An empirical analysis of terminological representation systems. DFKI Research Report RR-92-16, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, 1992. An extended and revised version of this paper has been published in Artificial Intelligence.

Heinsohn:1992:RRA

[HKNP92c] Jochen Heinsohn, Daniel Kudenko, Bernhard Nebel, and Hans-Jürgen Profitlich. RAT — representation of actions using terminological logics. In J. Heinsohn and B. Hollunder, editors, DFKI Workshop on Taxonomic Reasoning — Proceedings, number D-92-08 in DFKI Document. DFKI, Saarbrücken/Kaiserslautern, Germany, 1992.

Heinsohn:1994:EAT

[HKNP94] Jochen Heinsohn, Daniel Kudenko, Bernhard Nebel, and Hans-Jürgen Profitlich. An empirical analysis of terminological representation systems. Artificial Intelligence, 68(2):367–397, August 1994. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic). URL ftp:/

                                        /ftp.informatik.uni-ulm.
                                        de/pub/papers/ki/heinsohn-
                                        etal-aij-94.ps.

Isekenmeier:2002:EPC

[INW02] Guido Isekenmeier, Bernhard Nebel, and Thilo Weigel.

Evaluation of the performance of CS freiburg 1999 and CS freiburg 2000. Lecture Notes in Computer Science, 2377:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2377/23770393.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2377/23770393.
pdf.

Kobsa:1989:BWF

[KN89] Alfred Kobsa and Bernhard Nebel. Bericht über den Workshop on Formal Aspects of Semantic Networks. Künstliche Intelligenz, 89/2: 28, 1989. ISSN 0933-1875.

Koehler:1997:EPG

[KNHD97] Jana Koehler, Bernhard Nebel, Jörg Hoffmann, and Yannis Dimopoulos. Extending planning graphs to an ADL subset. Lecture Notes in Computer Science, 1348:273–285, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakemeyer:1994:FKRa

[LN94a] Gerhard Lakemeyer and Bernhard Nebel, editors. Foundations of Knowledge Representation, volume 810 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58107-3 (New York), 0-387-58107-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 355 pp. LCCN Q387 .F68 1994. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0810.htm;   http:
//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=810.

Lakemeyer:1994:FKR

[LN94b] Gerhard Lakemeyer and Bernhard Nebel. Foundations of knowledge representation: A guide to the volume. In Foundations of Knowledge Representation [LN94a], pages 1–12. ISBN 3-540-58107-3 (New York), 0-387-58107-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q387 .F68 1994. URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/lakemeyer-
nebel-94.ps.gz.

Li:2007:QSR

[LN07] Sanjiang Li and Bernhard Nebel. Qualitative spatial representation and reasoning: A hierarchical approach. The Computer Journal, 50 (4):391–402, July 2007. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/391;   http://comjnl.
oxfordjournals.org/cgi/
content/full/50/4/391;
http://comjnl.oxfordjournals.
org/cgi/reprint/50/4/391.

Marburger:1984:HAT

[MMN84] Heinz Marburger, Katharina Morik, and Bernhard Nebel. HAM-ANS on Tour: Bericht über eine USA-Reise. Memo ANS-24, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, 1984.

Marburger:1983:NDHa

[MN83a] Heinz Marburger and Bernhard Nebel. Natürlichsprachlicher Datenbankzugang mit HAM-ANS: Syntaktische Korrespondenz, natürlichsprachliche Quantifizierung und semantisches Modell des Diskursbereiches. In J. W. Schmidt, editor, Sprachen für Datenbanken, pages 26–

41. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983.

Marburger:1983:NDHb

[MN83b] Heinz Marburger and Bernhard Nebel. Natürlichsprachlicher Datenbankzugang mit HAM-ANS: Syntaktische Korrespondenz, natürlichsprachliche Quantifizierung und semantisches Modell des Diskursbereiches. Bericht ANS-22, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, September 1983. Published as [MN83a].

Mittelstein:1976:HPS

[MNPS76] Michael Mittelstein, Bernhard Nebel, Bernd Pretschner, and Peter Schefe. HASY — ein Programm zur syntaktischen Analyse natürlicher Sprache. Mitteilung IfI-HH-M36/76, Institut für Informatik, Universität Hamburg, Hamburg, Germany, October 1976.

Nebel:1991:CCT

[NB91] Bernhard Nebel and Christer Bäckström. On the computational complexity of temporal projection and some related problems. Research Report RR-91-34 (DFKI) and LiTH-IDA-R-91-34 (Univ. Linköping), German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, and Department of Computer and Information Science, Linköping University, Linköping, Sweden, October 1991.

Nebel:1992:CCT

[NB92] Bernhard Nebel and Christer Bäckström. On the computational complexity of temporal projection and plan validation. In AAAI-92. Proceedings, Tenth National Conference on Artificial Intelligence (San Jose, CA, 1992), pages 748–753. MIT Press, Cambridge, MA, USA, July 1992. An extended and revised version of this paper has been published in Artificial Intelligence.

Nebel:1993:RAT

[NB93] Bernhard Nebel and Hans-Jürgen Bürckert. Reasoning about temporal relations: A maximal tractable subclass of Allen’s interval algebra. DFKI Research Report RR-93-11, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, March 1993. A revised version was published in JACM.

Nebel:1994:CCT

[NB94a] Bernhard Nebel and Christer Bäckström. On the computational complexity of temporal projection, planning, and plan validation. Artificial Intelligence, 66 (1):125–160, 1994. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic). URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/nebel-
backstrom-aij-94.ps.

Nebel:1994:MQT

[NB94b] Bernhard Nebel and Hans-Jürgen Bürckert. Managing qualitative temporal information: Expressiveness vs. complexity. In K. von Luck and H. Marburger, editors, Management and Processing of Complex Data Structures, pages 104–117. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

Nebel:1994:RAT

[NB94c] Bernhard Nebel and Hans-Jürgen Bürckert. Reasoning about temporal relations: A maximal tractable subclass of Allen’s interval algebra. In Proceedings of the 12th National Conference of the American Association for Artificial Intelligence, pages 356–361. MIT Press, Cambridge, MA, USA, July 1994. Extended version published in Journal of the ACM.

Nebel:1995:RAT

[NB95] Bernhard Nebel and Hans-Jürgen Bürckert. Reasoning about temporal relations: A maximal tractable subclass of Allen’s interval algebra. Journal of the ACM, 42(1):43–66, January 1995. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/nebel-
burckert-jacm95.ps;   http:
//www.acm.org/pubs/toc/
Abstracts/0004-5411/200848.
html.

Nebel:1994:KAA

[NDF94] Bernhard Nebel and Leonie Dreschler-Fischer, editors. KI-94: advances in artificial intelligence: 18th German Annual Conference on Artificial Intelligence, Saarbrucken, Germany, September 18–23, 1994: proceedings, volume 861 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 3-540-58467-6 (Berlin), 0-387-58467-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .G466 1994. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0861.htm;   http:
//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=861.

Nebel:1997:IIF

[NDK97] Bernhard Nebel, Yannis Dimopoulos, and Jana Koehler. Ignoring irrelevant facts and operators in plan generation. Lecture Notes in Computer Science, 1348:338–350, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nebel:1980:WAA

[Neb80]  Bernhard Nebel.  Weit-

erführende Arbeiten am COPANET-Compiler: Realisierung der verdeckten Interrechner-Kommunikation im COPANET. Bericht IfI-HH-B-71/80, Institut für Informatik, Universität Hamburg, Hamburg, Germany, June 1980.

Nebel:1982:SHA

[Neb82] Bernhard Nebel. Der Systemrahmen von HAM-ANS. Memo ANS-7, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Hamburg, Germany, March 1982.

Nebel:1983:LSa

[Neb83a] Bernhard Nebel. Ist LISP eine ‘langsame’ Sprache? In B. Neumann, editor, GWAI-83. 7th German Workshop on Artificial Intelligence, pages 21–30. Springer-Verlag, Springer-Verlag, September 1983.

Nebel:1983:LSb

[Neb83b] Bernhard Nebel. Ist LISP eine ‘langsame’ Sprache? Memo ANS-13, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, February 1983. A short version has been published as [Neb83a].

Nebel:1984:UUL

[Neb84] Bernhard Nebel. ULM: Ein UCI-LISP — Lispmaschinen-LISP Übersetzungs-System.

Memo ANS-22, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, July 1984.

Nebel:1985:HWDa

[Neb85a] Bernhard Nebel. How well does a vanilla loop fit into a frame? Data and Knowledge Engineering, 1(2):181–194, 1985. CODEN DKENEW. ISSN 0169-023X (print), 1872-6933 (electronic).

Nebel:1985:HWDb

[Neb85b] Bernhard Nebel. How well does a vanilla loop fit into a frame? KIT Report 30, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, October 1985. Published as [Neb85a].

Nebel:1985:PWL

[Neb85c] Bernhard Nebel. Papiere zum Workshop “Logisches Programmieren und Lisp”. KIT Report 31, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, December 1985.

Nebel:1987:CCT

[Neb87a] Bernhard Nebel. Computational complexity of terminological reasoning in BACK. KIT Report 43, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, April 1987. Published as [Neb88].

Nebel:1987:TC

[Neb87b] Bernhard Nebel. On terminological cycles. KIT Report 58, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, November 1987.

Nebel:1988:CCT

[Neb88] Bernhard Nebel. Computational complexity of terminological reasoning in BACK. Artificial Intelligence, 34(3): 371–383, April 1988. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic). URL ftp:/

                                        /ftp.informatik.uni-ulm.
                                        de/pub/papers/ki/nebel-
                                        aij88.ps.gz.

Nebel:1989:KLA

[Neb89a] Bernhard Nebel. A knowledge level analysis of belief revision. In R. Brachman, H. J. Levesque, and R. Reiter, editors, Proceedings of the First International Conference on Principles of Knowledge Representation and Reasoning (Toronto, ON, 1989), Morgan Kaufmann Ser. Represent. Reason., pages 301– 311. Morgan Kaufmann, San Mateo, May 1989.

Nebel:1989:KLAb

[Neb89b] Bernhard Nebel. A knowledge level analysis of belief revision. IWBS Report 69, IWBS, IBM Germany, Stuttgart, Germany, 1989.

Nebel:1989:TRI

[Neb89c] Bernhard Nebel. Terminological reasoning is inherently intractable. IWBS Report 82, IWBS, IBM Germany, Stuttgart, Germany, 1989.

Nebel:1989:WHH

[Neb89d] Bernhard Nebel. What is hybrid in hybrid representation systems? In F. Gardin, G. Mauri, and M. G. Filippini, editors, Proceedings of the International Symposium on Computational Intelligence ’89, pages 217–228. North-Holland, Amsterdam, The Netherlands, September 1989.

Nebel:1990:RRH

[Neb90a] Bernhard Nebel. Reasoning and Revision in Hybrid Representation Systems, volume 422 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-52443-6 (Berlin), 0-387-52443-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 270 pp. LCCN Q335 .N43 1990. URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/reprint.
ps.gz;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/tocs/t0422.htm;   http:
                                        //www.springerlink.com/
                                        openurl.asp?genre=issue&
                                        issn=0302-9743&volume=422.

Nebel:1990:TRI

[Neb90b] Bernhard Nebel. Terminological reasoning is inherently intractable. Artificial Intelligence, 43(2):235–249, 1990. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic). URL ftp:/

                                        /ftp.informatik.uni-ulm.
                                        de/pub/papers/ki/nebel-
                                        aij90.ps.gz.

Nebel:1991:BRD

[Neb91a] Bernhard Nebel. Belief revision and default reasoning: syntax-based approaches. In Principles of knowledge representation and reasoning (Cambridge, MA, 1991), Morgan Kaufmann Ser. Represent. Reason., pages 417– 428. Morgan Kaufmann, San Mateo, CA, 1991.

Nebel:1991:BRDb

[Neb91b] Bernhard Nebel. Belief revision and default reasoning: Syntax-based approaches. In J. A. Allen, R. Fikes, and E. Sandewall, editors, Principles of Knowledge Representation and Reasoning: Proceedings of the 2nd International Conference, pages 417– 428. Morgan Kaufmann, San Mateo, April 1991.

Nebel:1991:BRDa

[Neb91c]  Bernhard Nebel.  Belief

revision and default reasoning: Syntax-based approaches. DFKI Research Report RR-91-11, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, 1991.

Nebel:1991:CMI

[Neb91d] Bernhard Nebel. The complexity of multiple inheritance in complex object data models. In Workshop Notes of the IJCAI’91 Workshop on “Objects and AI”. IJCAI, Sydney, Australia, 1991.

Nebel:1991:TCS

[Neb91e] Bernhard Nebel. Terminological cycles: Semantics and computational properties. In John F. Sowa, editor, Principles of Semantic Networks, pages 331–362. Morgan Kaufmann, San Mateo, 1991.

Nebel:1992:CCK

[Neb92a]  Bernhard Nebel.  Com-

putational complexity and KR&R: Is polynomial time all that matters? In Workshop Notes of the AAAI’92 Workshop on “Tractable Reasoning”, pages 126–129. AAAI, San Jose, CA, 1992.

Nebel:1992:SBA

[Neb92b] Bernhard Nebel. Syntaxbased approaches to belief revision. In P. Gärdenfors, editor, Belief Revision, volume 29 of Cambridge Tracts in Theoretical Computer Science, pages 52–88. Cambridge University Press, Cambridge, UK, 1992. URL

ftp://ftp.informatik.uni-
ulm.de/pub/papers/ki/nebel-
br92.ps.gz.

Nebel:1992:TLR

[Neb92c] Bernhard Nebel. Terminological logics and representation systems. In J. Heinsohn and B. Hollunder, editors, DFKI Workshop on Taxonomic Reasoning — Proceedings, number D-92-08 in DFKI Document. DFKI, Saarbrücken/Kaiserslautern, Germany, 1992.

Nebel:1994:BRO

[Neb94] Bernhard Nebel. Base revision operations and schemes: Semantics, representation, and complexity. In Proceedings of the 11th European Conference on Artificial Intelligence, pages 341–345. Wiley, New York, NY, USA, August 1994. URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/nebel-
ecai-94.ps.

Nebel:1995:AIC

[Neb95a]  Bernhard Nebel.  Artifi-

cial intelligence: A computational perspective. In G. Brewka, editor, Essentials in Knowledge Representation, Studies in Logic, Language and Information. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1995. URL ftp:/

/ftp.informatik.uni-ulm.
de/pub/papers/ki/nebel-
lli93-book.ps. To appear.

Nebel:1995:BRO

[Neb95b] Bernhard Nebel. Base revision operations and schemes: Semantics, representation, and complexity. In G. della Riccia, R. Kruse, and R. Viertl, editors, Mathematical and statistical methods in artificial intelligence (Udine, 1994), volume 363 of CISM Courses and Lectures, pages 157–170. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. Revised version of [Neb94].

Nebel:1995:CPQ

[Neb95c] Bernhard Nebel. Computational properties of qualitative spatial reasoning: First results. In I. Wachsmuth, C.- R. Rollinger, and W. Brauer, editors, KI-95: Advances in artificial intelligence (Bielefeld, 1995), volume 981 of Lecture Notes in Comput. Sci., pages 233–244. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995.

Nebel:1995:KKI

[Neb95d]  Bernhard Nebel.  Kom-

plexitätsanalysen in der Künstlichen Intelligenz. Künstliche Intelligenz, 95/2:6–14, 1995. ISSN 0933-1875. URL ftp:/

                                        /ftp.informatik.uni-ulm.
                                        de/pub/papers/ki/nebel94.
                                        ps.

Nebel:1996:AIC

[Neb96] Bernhard Nebel. Artificial intelligence: a computational perspective. In Principles of knowledge representation, Stud. Logic Lang. Inform., pages 237–266. CSLI Publ., Stanford, CA, 1996.

Nebel:1997:SHQ

[Neb97] Bernhard Nebel. Solving hard qualitative temporal reasoning problems: evaluating the efficiency of using the ORD–Horn class. Constraints, 1(3):175–190, 1997. CODEN CNSTFT. ISSN 1383-7133.

Nebel:1998:HHI

[Neb98] Bernhard Nebel. How hard is it to revise a belief base? In Belief change, volume 3 of Handb. Defeasible Reason. Uncertain. Manag. Syst., pages 77–145. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998.

Nebel:2000:CEP

[Neb00] Bernhard Nebel. On the compilability and expressive power of propositional planning formalisms. J. Artificial Intelligence Res., 12:271–315, 2000. ISSN 1076-9757.

Nebel:2001:CPR

[Neb01] Bernhard Nebel. Cooperating physical robots: A lesson in playing robotic soccer. Lecture Notes in Computer Science, 2086:404–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2086/20860404.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2086/20860404.
pdf.

Nebel:1992:PMV

[NK92] Bernhard Nebel and Jana Koehler. Plan modification versus plan generation: A complexity-theoretic perspective. DFKI Research Report RR-92-48, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, November 1992.

Nebel:1993:PMV

[NK93a] Bernhard Nebel and Jana Koehler. Plan modification versus plan generation: A complexity-theoretic perspective. In Proceedings of the 13th International Joint Conference on Artificial Intelligence, pages 1436–1441. Morgan Kaufmann, San Mateo, August 1993.

Nebel:1993:PRV

[NK93b] Bernhard Nebel and Jana Koehler. Plan reuse versus plan generation: A theoretical and empirical analysis. DFKI Research Report RR-93-33, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, 1993. A revised and extended version was published in Artificial Intelligence, Special Issue on Planning and Scheduling.

Nebel:1995:PRV

[NK95] Bernhard Nebel and Jana Koehler. Plan reuse versus plan generation: A theoretical and empirical analysis. Artificial Intelligence, 76 ((1-2)):427–454, 1995. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic). URL ftp:/

                                        /ftp.informatik.uni-ulm.
                                        de/pub/papers/ki/nebel-
                                        koehler-aij95.ps.

Nebel:1982:NSHa

[NM82a] Bernhard Nebel and Heinz Marburger. Das natürlichsprachliche System HAM-ANS: Intelligenter Zugriff auf heterogene Wissens- und Datenbasen. In J. Nehmer, editor, GI-

12. Jahrestagung, pages 392–

402. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982.

Nebel:1982:NSHb

[NM82b] Bernhard Nebel and Heinz Marburger. Das natürlichsprachliche System HAM-ANS: Intelligenter Zugriff auf heterogene Wissens- und Datenbasen. Bericht ANS-7, Forschungsstelle für Informationswissenschaft und Künstliche Intelligenz, Universität Hamburg, Hamburg, Germany, 1982. Published as [NM82a].

Nebel:1976:EDP

[NP76] Bernhard Nebel and Bernd Pretschner. Erweiterung des DECSystem-10 PASCAL-Compilers um eine Möglichkeit zur Erzeugung eines Post-Mortem-Dump. Mitteilung IfI-HH-M34/76, Institut für Informatik, Universität Hamburg, Hamburg, Germany, June 1976.

Nebel:1991:TRI

[NP91a] Bernhard Nebel and Christof Peltason. Terminological reasoning and information management. In Information systems and artificial intelligence: integration aspects (Ulm, 1990), volume 474 of Lecture Notes in Comput. Sci., pages 181–212. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Nebel:1991:TRIb

[NP91b] Bernhard Nebel and Christof Peltason. Terminological reasoning and information management. In D. Karagianis, editor, Information Systems and Artificial Intelligence: Integration Aspects, pages 181–212. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Nebel:1991:TRIa

[NP91c] Bernhard Nebel and Christof Peltason. Terminological reasoning and information management. Kit report, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, 1991.

Nebel:1991:IWTa

[NPvL91a] Bernhard Nebel, Christof Peltason, and Kai von Luck. International workshop on terminological logics. DFKI Document D-91-13, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, 1991. Also published as KIT Report 89, TU Berlin, and IWBS Report 184, IBM Germany, Stuttgart.

Nebel:1991:IWTb

[NPvL91b] Bernhard Nebel, Christof Peltason, and Kai von Luck. International workshop on terminological logics — seminar report. Dagstuhl-Seminar-Report 12 (9119), Schloß Dagstuhl, Germany, 1991.

Nebel:1985:NGK

[NS85] Bernhard Nebel and Norman K. Sondheimer. NIGEL gets to know logic: An experiment in natural language generation taking a logical, knowledge-based view. KIT Report 36, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, July 1985. Published as [NS86].

Nebel:1986:NGK

[NS86] Bernhard Nebel and Norman K. Sondheimer. NIGEL gets to know logic: An experiment in natural language generation taking a logical, knowledge-based view. In C.- R. Rollinger and W. Horn, editors, GWAI-86 und 2. Österreichische Artificial-Intelligence-Tagung, pages 75–86. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1986.

Nebel:1989:RRA

[NS89] Bernhard Nebel and Gert Smolka. Representation and reasoning with attributive descriptions. IWBS Report 81, IWBS, IBM Germany, Stuttgart, Germany, 1989.

Nebel:1990:RRA

[NS90] Bernhard Nebel and Gert Smolka. Representation and reasoning with attributive descriptions. In K.-H. Bläsius, U. Hedtstück, and C.-R. Rollinger, editors, Sorts and types in artificial intelligence (Eringerfeld, 1989), volume 418 of Lecture Notes in Artificial Intelligence, pages 112– 139. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990.

Nebel:1991:ADFa

[NS91a] Bernhard Nebel and Gert Smolka. Attributive description formalisms … and the rest of the world. In Otthein Herzog and Claus-Rainer Rollinger, editors, Text Understanding in LILOG, volume 546 of Lecture Notes in Artificial Intelligence, pages 439–452. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991.

Nebel:1991:ADFb

[NS91b] Bernhard Nebel and Gert Smolka. Attributive description formalisms … and the rest of the world. DFKI Research Report RR-91-15, German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany, 1991.

Nebel:1992:PKR

[NSR92] B. Nebel, W. Swartout, and C. Rich, editors. Principles of Knowledge Representation and Reasoning: Proceedings of the 3rd International Conference, The Morgan Kaufmann series in representation and reasoning. Morgan Kaufmann, San Mateo, October 1992. ISBN 1-55860-262-3. ISSN 1046-9567. LCCN Q387 .P76 1992.

Nebel:1987:IIBa

[NvL87a] Bernhard Nebel and Kai von Luck. Issues of integration and balancing in hybrid knowledge representation systems. In K. Morik, editor, GWAI-87. 11th German Workshop on Artificial Intelligence, pages 114–123. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1987.

Nebel:1987:IIBb

[NvL87b] Bernhard Nebel and Kai von Luck. Issues of integration and balancing in hybrid knowledge representation systems. KIT Report 46, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, July 1987. Published as [NvL87a].

Nebel:1988:HRB

[NvL88] Bernhard Nebel and Kai von Luck. Hybrid reasoning in BACK. In Z. W. Ras and L. Saitta, editors, Proceedings of the Third International Symposium on Methodologies for Intelligent systems, pages 260–269. North-Holland, Amsterdam, The Netherlands, October 1988.

Owsnicki-Klewe:1993:WLE

[OKvLN93] Bernd Owsnicki-Klewe, Kai von Luck, and Bernhard Nebel. Wissensrepräsentation und Logik — Eine Einführung. In G. Görz, editor, Einführung in die Künstliche Intelligenz, chapter 1.1, pages 3– 54. Addison-Wesley Publishing Company, Reading, MA, USA, 1993.

Profitlich:1991:CAT

[PHKN91] Hans-Jürgen Profitlich, Jochen Heinsohn, Daniel Kudenko, and Bernhard Nebel. A comparative analysis of terminological representation systems. In Working Notes of the AAAI Spring Symposium 1991 on “Implemented Knowledge Representation Systems”, pages 347– 360. AAAI, Stanford, CA, 1991. Preliminary version of [HKNP94].

Padgham:1993:CCN

[PN93] Lin Padgham and Bernhard Nebel. Combining classification and nonmonotonic inheritance reasoning: A first step. In J. Komorowski and Z. W. Ras, editors, Proceedings of the Seventh International Symposium on Methodologies for Intelligent systems, Trondheim, 1993, volume 689 of Lecture Notes in Comput. Sci., pages 132– 141. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., June 1993.

Patel-Schneider:1990:TSL

[PSOKK^{+}90] Peter F. Patel-Schneider, Bernd Owsnicki-Klewe, Alfred Kobsa, Nocola Guarino, Robert MacGregor, William S. Mark, Deborah McGuinness, Bernhard Nebel, Albrecht Schmiedel, and John Yen. Term subsumption languages in knowledge representation. The AI Magazine, 11(2):16–23, 1990. CODEN AIMAEK. ISSN 0738-4602.

Peltason:1987:UGB

[PvLNS87] Christof Peltason, Kai von Luck, Bernhard Nebel, and Albrecht Schmiedel. The user’s guide to the BACK system. KIT Report 42, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, January 1987.

Renz:1998:SRT

[RN98] Jochen Renz and Bernhard Nebel. Spatial reasoning with topological information. Lecture Notes in Computer Science, 1404:351–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1404/14040351.
htm;   http://link.springer-
ny.com/link/service/series/

                                        0558/papers/1404/14040351.
                                        pdf.

Renz:1999:CQS

[RN99] Jochen Renz and Bernhard Nebel. On the complexity of qualitative spatial reasoning: a maximal tractable fragment of the region connection calculus. Artificial Intelligence, 108(1-2):69–123, 1999. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Renz:2001:EMQ

[RN01] Jochen Renz and Bernhard Nebel. Efficient methods for qualitative spatial reasoning. J. Artificial Intelligence Res., 15:289–318, 2001. ISSN 1076-9757.

Renz:2007:QSR

[RN07] Jochen Renz and Bernhard Nebel. Qualitative spatial reasoning using constraint calculi. In Handbook of spatial logics, pages 161–215. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007.

Sondheimer:1986:LFKa

[SN86a] Norman K. Sondheimer and Bernhard Nebel. A logicalform and knowledge-base design for natural language generation. In Proceedings of the 5th National Conference of the American Association for Artificial Intelligence, pages 612–618. Morgan Kaufmann, Philadelphia, PA, August 1986.

Sondheimer:1986:LFKb

[SN86b] Norman K. Sondheimer and Bernhard Nebel. A logicalform and knowledge-base design for natural language generation. Reprint Series ISI/RS-86-169, University of Southern California, Information Science Institute, Marina del Rey, CA, November 1986. Reprint of [SN86a].

Scivos:2001:DCD

[SN01] Alexander Scivos and Bernhard Nebel. Double-crossing: Decidability and computational complexity of a qualitative calculus for navigation. Lecture Notes in Computer Science, 2205:431–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2205/22050431.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2205/22050431.
pdf.

Schmiedel:1986:PCS

[SvLNP86] Albrecht Schmiedel, Kai von Luck, Bernhard Nebel, and Christof Peltason. ‘Bitter Pills’: A case study in knowledge representation. KIT Report 39, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, August 1986.

Thiebaux:2005:DPA

[THN05] Sylvie Thiébaux, Jörg Hoffmann, and Bernhard Nebel. In defense of PDDL axioms. Artificial Intelligence, 168(1-2):38–69, 2005. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Luck:1985:BS

[vLNPS85a] Kai von Luck, Bernhard Nebel, Christof Peltason, and Albrecht Schmiedel. The BACK system. KIT Report 28, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, August 1985.

Luck:1985:BCI

[vLNPS85b] Kai von Luck, Bernhard Nebel, Christof Peltason, and Albrecht Schmiedel. BACK to consistency and incompleteness. In H. Stoyan, editor, GWAI-85. 9th German Workshop on Artificial Intelligence, pages 245– 257. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1985.

Luck:1987:ABS

[vLNPS87] Kai von Luck, Bernhard Nebel, Christof Peltason, and Albrecht Schmiedel. The anatomy of the BACK system. KIT Report 41, Department of Computer Science, Technische Universität Berlin, Berlin, Germany, January 1987.

Luck:1988:SAK

[vLNS88] Kai von Luck, Bernhard Nebel, and Hans-Jochen Schneider. Some aspects of knowledge-base management systems. In G. Rahmstorf, editor, Wissensrepräsentation in Expertensystemen, pages 146–157. Springer-Verlag, Berlin, Heidelberg, New York, 1988.

Weigel:2001:CFD

[WAD^{+}01] Thilo Weigel, Willi Auerbach, Markus Dietl, Burkhard Dümler, Jens-Steffen Gutmann, Kornel Marko, Klaus Müller, Bernhard Nebel, Boris Szerbakowski, and Maximilian Thiel. CS freiburg: Doing the right thing in a group. Lecture Notes in Computer Science, 2019:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2019/20190052.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2019/20190052.
pdf.

Weigel:2002:CF

[WKD^{+}02] Thilo Weigel, Alexander Kleiner, Florian Diesch, Markus Dietl, Jens-Steffen Gutmann, Bernhard Nebel, Patrick Stiegeler, and Boris Szerbakowski. CS freiburg 2001. Lecture Notes in Computer Science, 2377:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2377/23770026.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2377/23770026.
                                        pdf.